Find the missing value. the principal is $500, the time is 7 years, the simple interest is $140.00, i have to find the rate.
1
Expert's answer
2012-12-03T09:18:30-0500
The simple interest formula is:
I = P*R*T, where
I - interest, P - principal, R - rate, T - time in years of the loan.
Then, R = I/(P*T) =140/(500*7) = 0.04
The rate is 4%